topic:- integrated Enterprise Systems

* no hand writing

*the answer must be unique not copied “plagiarized “


1- SAM creates a production order for the requested 12 cars. For each car, raw materials are expected to cost $750, wheel assemblies (semi finished goods) $430. The actual time (labor) required to build one car is 49 minutes and hourly rate is $50


2- don’t forget to do the Calculate the estimated cost of the production order and provide the financial impact using" T" account
Answer #1

Given for a car the raw materials cost $750 and wheel assemblies cost $430 which totals to $1180.

This equals to 12*1180 = $14160

And it takes 49*12 minutes = 588 min = 9.8 hours for 12 cars which costs 9.8*50 = $490

So it costs a total of $14160+490 = $14650 for 12 cars